True/False . Why?
Suppose the country of Macecon produces only two final goods, Theorems (Ts) and Lemmas (Ls). In 2018 it produced 10 Ts at a price of $50 and 50 Ls at a price of $10. In 2019 It produced 9 Ts at a price of $60 and 60 Ls at a price of $11. Using 2018 as the base year, the inflation rate in the GDP deflator between 2018 and 2019 was 5%.
1
Expert's answer
2020-04-20T10:31:59-0400

False

GDP deflator = "\frac{nominal GDP}{real GDP}\times100"

Nominal GDP 2018 = (10 x 50) + (50 x 10) = 1000

Nominal GDP 2019 = (9 x 60) + (60 x 11) = 1210

Real GDP 2018 =(10 x 50) + (50 x 10) = 1000

Real GDP 2019 = (9 x 50) + (60 x 10) = 1050

GDP deflator 2019 = (1210/1050) x 100 = 115.24%

Inflation rate = 115.24 - 100 = 15.24%

Therefore, inflation rate is 15.24% and not 5%